# On the $\ell^{2}$-Betti numbers and algebraic fibring of the (outer) automorphism group of a right-angled Artin groups

### v2

@@ 3.2 Transvection subgroup of $\operatorname{Out}(A_{\Gamma})$
We define a directed graph $\Lambda_{\Gamma}$, called transvection graph, with vertex set $\{V_{1},\dots,V_{N}\}$ and such that there is a directed edge from $V_{i}$ to $V_{j}$ if $v\leq w$ for some $v\in V_{i}$ and $w\in V_{j}$. There is a loop from $V_{i}$ to itself if and only if $\left|V_{i}\right|\geq 2$. In this setting, the image of a transvection $\tau_{v_{i}}^{v_{j}}$ under $\rho$ is the elementary matrix $\mathrm{E}_{i,j}$ with $1$’s in the diagonal and in the $(i,j)$-entry and zeroes elsewhere. Thus, the matrices in $Q$ are block lower triangular, with diagonal blocks in $\mathrm{SL}_{\left|V_{i}\right|}(\mathbb{Z})$ and non-trivial off-diagonal blocks $(i,j)$ permitted precisely hen there is an edge $V_{i}\to V_{j}$ in $\Lambda_{\Gamma}$.
@@
    There exists two singleton equivalence classes $u]=\{u\}$ and $v]=\{v\}$ with $u\leq v$, such that there is no vertex $w\neq u,v$ satisfying $u\leq w\leq v$.
@@
If property (P2) holds then there exists $v_{i}\leq v_{j}$ such that there is no $v_{r}\neq v_{i},v_{j}$ with $v_{i}\leq v_{r}\leq v_{j}$. Moreover, $v_{i}]=\{v_{i}\}$ and $v_{j}]=\{v_{j}\}$. Consider the set:
@@
and define the map $f:Q\to\mathbb{Z}$ by $f(E_{i,j})=1$ and $f(E_{k,l})=0$ if $E_{k,l}\in\mathcal{E}$. This is a well-defined group homomorphism because of the presentation of $Q$ from Theorem 3.10. We claim that $\ker(f)=\langle\mathcal{E}\rangle$. As $\ker(f)$ is normally generated by $\mathcal{E}$, it is enough to prove that for every $E_{k,l}\in\mathcal{E}$ there exists $R\in\langle\mathcal{E}\rangle$ such that $E_{i,j}E_{k,l}=RE_{i,j}$. If $i\neq l$ and $j\neq k$ then $E_{i,j}$ and $E_{k,l}$ commute so the claim follows. If $i=l$ and $j\neq k$ then $E_{k,j}\in\mathcal{E}$ and $E_{i,j}E_{k,i}=E_{k,i}E_{k,j}E_{i,j}$ and if $j=k$ and $i\neq l$ then $E_{i,l}\in\mathcal{E}$ and $E_{i,j}E_{j,l}=E_{j,l}(E_{i,l})^{-1}E_{i,j}$. We deduce that $\ker(f)$ is finitely generated.
@@
3.3 Fibring of the automorphism group oa a direct product
A group $G$ is said to be indicable if there is a non-trivial homomorphism $\varphi:G\to\mathbb{Z}$. The following result provides a method for constructing groups that algebraically fibre.
Proposition 3.16.
Let $G$ and $H$ be finitely generated indicable groups. Then, $G\times H$ algebraically fibres.
Proof.
Follows directly from [18, Theorem 1]. ∎
In the context of automorphism groups this result yields the following:
Proposition 3.17.
Let $G$ and $H$ be directly indecomposable, centreless groups. Then:
  1. 1.
    If $\operatorname{Aut}(G)$ and $\operatorname{Aut}(H)$ are finitely generated, virtually indicable groups, then $\operatorname{Aut}(G\times H)$ virtually algebraically fibres.
  2. 2.
    If $\operatorname{Out}(G)$ and $\operatorname{Out}(H)$ are finitely generated, virtually indicable groups, then $\operatorname{Out}(G\times H)$ virtually algebraically fibres.
Proof.
By Proposition 2.21 $\operatorname{Aut}(G)\times\operatorname{Aut}(H)$ is a finite index subgroup of $\operatorname{Aut}(G\times H)$. Since $\operatorname{Aut}(G)\times\operatorname{Aut}(H)$ virtually algebraic fibres by Proposition 3.16, the result follows. The same argument applies to $\operatorname{Out}(G\times H)$. ∎
For a RAAG $A_{\Gamma}$, there are conditions under which either $\operatorname{Aut}(A_{\Gamma})$ or $\operatorname{Out}(A_{\Gamma})$ is virtually indicable, allowing us to apply Proposition 3.17. We first introduce a relevant notion:
Definition 3.18.
Let $v\in V(\Gamma)$ and $C$ a connected component of $\Gamma\setminus\operatorname{st}(v)$. We say that the partial conjugation $\pi_{C}^{v}$ is virtually obtained from dominated components if there exists $n\neq 0$ such that in $\operatorname{Out}(A_{\Gamma})$:
$(\pi^{v}_{C})^{n}=(\pi^{v}_{D_{1}})^{\epsilon_{1}}\cdots(\pi^{v}_{D_{m}})^{\epsilon_{m}}$
where each $D_{i}$ is a component of $\Gamma\setminus\operatorname{st}(y_{i})$ for some $y_{i}\lneq v$ and $\epsilon_{i}\in\{\pm 1\}$.
We now recall several conditions that ensure virtual indicability of $\operatorname{Aut}(A_{\Gamma})$ or $\operatorname{Out}(A_{\Gamma})$.
Theorem 3.19.
Let $A_{\Gamma}$ be a RAAG. Then:
  1. 1.
    If there exists $u,v\in V(\Gamma)$ with $u\lneq v$ and there is no $w\in V(\Gamma)$ such that $u\lneq w\lneq v$ then $\operatorname{Aut}(A_{\Gamma})$ is virtually indicable. [[2] Corollary 1.4, Aramayona-Martinez]
  2. 2.
    If there exists $w\in V(\Gamma)$ such that there is no $v\in V(\Gamma)$ with $v\lneq w$, then $\operatorname{Aut}(A_{\Gamma})$ is virtually indicable. [[2] Theorem 1.6, Aramayona-Martinez]
  3. 3.
    If there exists $v\in V(\Gamma)$ and a component $C$ of $\Gamma\setminus\operatorname{st}(v)$ such that the partial conjugation $\pi_{C}^{v}$ is not virtually obtained from dominated components, then $\operatorname{Out}(A_{\Gamma})$ is virtually indicable. [[29] Theorem 3, Sale]
Condition (3) can be difficult to verify in practice. However, the following weaker condition is easier to check and also implies virtual indicability:
  • 3’.
    There exists $w\in V(\Gamma)$ such $\Gamma\setminus\operatorname{st}(w)$ is not connected and such that there is no $v\in V(\Gamma)$ with $v\lneq w$. [[2] Theorem 1.6, Aramayona-Martinez]
We will illustrate this method with explicit examples in Section 5.3.
@@ 4.1 Transvection subgroup of $\operatorname{Out}(A_{\Gamma})$
  • If there are partial conjugations but no transvections, then $\mathrm{PSO}(A_{\Gamma})$ is a finite index subgroup of $\operatorname{Out}(A_{\Gamma})$. By Proposition 3.7, we have that $\mathrm{PSO}(A_{\Gamma})$ virtually fibres if and only if $\mathrm{PSO}(A_{\Gamma})$ is not isomorphic to a RAAG $A_{\Theta}$ with $\Theta$ disconnected. Since virtual fibring implies vanishing of the first $\ell^{2}$-Betti number, it follows that $\beta_{1}^{(2)}(\operatorname{Out}(A_{\Gamma}))>0$ if and only if $\mathrm{PSO}(A_{\Gamma})$ is isomorphic to a RAAG $A_{\Theta}$ with $\Theta$ a disconnected graph.∎
4.3 $\ell^{2}$-Betti numbers of $\operatorname{Out}(A_{\Gamma})$ with $\Gamma$ disconnected
In this section we will prove Proposition 1.3. To do so, we need to introduce the relative outer automorphism group of a RAAG. A complete description of those groups was given by Day and Wade in [15]. We begin by recalling the relevant definitions and the main results from [15] that we will use.
Definition 4.4.
Let $A_{\Gamma}$ be a RAAG and let $A_{\Delta}\leq A_{\Gamma}$ be an special subgroup.
  • An outer automorphism $\Phi\in\operatorname{Out}(A_{\Gamma})$ is said to preserve $A_{\Delta}$ if there exists a representative $\phi\in\Phi$ that restricts to an automorphism of $A_{\Delta}$.
  • An outer automorphism $\Phi\in\operatorname{Out}(A_{\Gamma})$ is said to act trivially on $A_{\Delta}$ if there exists a representative $\phi\in\Phi$ acting as the identity on $A_{\Delta}$.
Let $\mathcal{G}$ and $\mathcal{H}$ be collections of special subgroups of $A_{\Gamma}$. Define the relative outer automorphism group $\operatorname{Out}(A_{\Gamma},\mathcal{G},\mathcal{H}^{t})$ as the subgroup of $\operatorname{Out}(A_{\Gamma})$ consisting of all outer automorphisms that preserve each element of $\mathcal{G}$ and act trivially on each element of $\mathcal{H}$. We also define the finite index subgroup:
$\operatorname{Out}^{0}(A_{\Gamma},\mathcal{G},\mathcal{H}^{t})=\operatorname{Out}(A_{\Gamma},\mathcal{G},\mathcal{H}^{t})\cap\operatorname{Out}^{0}(A_{\Gamma})$
We say that $\mathcal{G}$ is saturated with respect to $(\mathcal{G},\mathcal{H})$ if it contains every proper special subgroup of $A_{\Gamma}$ that is preserved under $\operatorname{Out}^{0}(A_{\Gamma},\mathcal{G},\mathcal{H}^{t})$. Given a special subgroup $A_{\Delta}$ we denote:
$\mathcal{G}_{\Delta}=\{A_{\Delta\cap\Theta}\mid A_{\Theta}\in\mathcal{G}\}\setminus\{A_{\Delta}\},~\mathcal{H}_{\Delta}=\{A_{\Delta\cap\Theta}\mid A_{\Theta}\in\mathcal{H}\}\setminus\{A_{\Delta}\}$
Theorem 4.5 ([15] Theorem E).
Let $\mathcal{G}$ and $\mathcal{H}$ be collections of special subgroups of $A_{\Gamma}$. Let $A_{\Delta}\in\mathcal{G}$ and suppose that $\mathcal{G}$ is saturated with respect to $(\mathcal{G},\mathcal{H})$. Then, there is a short exact sequence:
$1\to\operatorname{Out}^{0}(A_{\Gamma},\mathcal{G},(\mathcal{H}\cup\{A_{\Delta}\})^{t})\to\operatorname{Out}^{0}(A_{\Gamma},\mathcal{G},\mathcal{H}^{t})\to\operatorname{Out}^{0}(A_{\Delta},\mathcal{G}_{\Delta},\mathcal{H}_{\Delta}^{t})\to 1$
There is a canonical generating set for $\operatorname{Out}^{0}(A_{\Gamma},\mathcal{G},\mathcal{H}^{t})$. Indeed, in [15] Day and Wade proved that $\operatorname{Out}^{0}(A_{\Gamma},\mathcal{G},\mathcal{H}^{t})$ is generated by the subset of inversions, transvections and extended partial conjugations it contains.
Lemma 4.6.
Let $\Gamma$ be a disconnected simplicial graph and $\Delta$ a connected component of $\Gamma$ with at leat two vertices. Let $\mathcal{G}$ be the set of those special subgroups of $A_{\Gamma}$ which are preserved by $\operatorname{Out}^{0}(A_{\Gamma})$. Then, $A_{\Delta}\in\mathcal{G}$ and $\pi_{\Delta}^{v}]\in\mathrm{Z}(\operatorname{Out}^{0}(A_{\Gamma},\mathcal{G},\{A_{\Delta}\}^{t}))$ for any $v\in V(\Gamma)$.
Proof.
It is straightforward to check that every standard generator of $\operatorname{Aut}^{0}(A_{\Gamma})$ restricts to a well-defined automorphism of $\operatorname{Aut}^{0}(A_{\Delta})$. Hence, $A_{\Delta}$ is preserved by $\operatorname{Aut}^{0}(A_{\Delta})$ and so $A_{\Delta}\in\mathcal{G}$. Consider the standard generators of $\operatorname{Out}^{0}(A_{\Gamma},\mathcal{G},\{A_{\Delta}\}^{t})$: cosets of inversions $\iota_{w}$ with $w\notin V(\Delta)$, transvections $\tau^{u}_{w}$ with $w\notin V(\Delta)$ and partial conjugations $\pi_{C}^{w}$ with either $C\cap\Delta=\emptyset$ or $\Delta\subset C$. However, any partial conjugation $\pi_{C}^{w}$ satisfies the identity $\pi_{C}^{w}]=[\pi^{w}_{\Gamma\setminus C}]^{-1}$ in $\operatorname{Out}(A_{\Gamma})$. Thus, without loss of generality, we may choose the partial conjugations $\pi_{C}^{w}$ such that $C\cap\Delta=\emptyset$. In this case, the support of each standard generator is disjoint from $\Delta$, and hence these generators commute with $\pi_{\Delta}^{v}$. It follows that $\pi_{\Delta}^{v}\in\mathrm{Z}(\operatorname{Out}^{0}(A_{\Gamma},\mathcal{G},(\mathcal{H}\cup\{A_{\Gamma_{1}}\})^{t}))$ and the claim follows. ∎
Proposition 4.7.
Let $A_{\Gamma}=A_{\Gamma_{1}}\ast\cdots\ast A_{\Gamma_{n}}$ be the maximal free product decomposition of a RAAG. Suppose one of the following conditions holds:
  1. 1.
    $\Gamma_{i}$ is not a complete graph for some $i$.
  2. 2.
    $n\geq 3$ and $\Gamma_{i}$ has at least $2$ vertices for some $i$.
Then, all $\ell^{2}$-Betti numbers of $\operatorname{Out}(A_{\Gamma})$ vanish.
Proof.
Let $\mathcal{G}$ be the set of those special subgroups of $A_{\Gamma}$ which are preserved by $\operatorname{Out}^{0}(A_{\Gamma})$.
By Lemma 4.6 we have $A_{\Gamma_{1}}\in\mathcal{G}$. Hence, we can consider the short exact sequence from Theorem 4.5:
$1\to\operatorname{Out}^{0}(A_{\Gamma},\mathcal{G},\{A_{\Gamma_{1}}\}^{t})\to\operatorname{Out}^{0}(A_{\Gamma},\mathcal{G})\to\operatorname{Out}^{0}(A_{\Gamma_{1}},\mathcal{G}_{\Gamma_{1}})\to 1$
To prove that the $\ell^{2}$-Betti numbers of $\operatorname{Out}^{0}(A_{\Gamma})$ vanish, it suffices to show that the same holds for $\operatorname{Out}^{0}(A_{\Gamma},\mathcal{G},\{A_{\Gamma_{1}}\}^{t})$. Since the centre of a group is a normal abelian subgroup, having an infinite centre implies that all $\ell^{2}$-Betti numbers vanish. Thus, we just need to show that the centre of $\operatorname{Out}^{0}(A_{\Gamma},\mathcal{G},\{A_{\Gamma_{1}}\}^{t})$ is non-trivial.
From Lemma 4.6 we have $\pi_{\Gamma_{i}}^{v}]\in\mathrm{Z}(\operatorname{Out}^{0}(A_{\Gamma},\mathcal{G},\{A_{\Gamma_{1}}\}^{t}))$ for all vertices $v\in V(\Gamma)$. For 1. we can choose $v\in V(\Gamma_{1})$ such that $\operatorname{st}(v)\neq\Gamma_{1}$, which exist since $\Gamma_{1}$ is non complete. In the second case we can consider $v\in V(\Gamma_{2})$. In both cases the centre contains an infinite cyclic subgroup generated by $\pi_{\Gamma_{1}}^{v}]$ so it is infinite. ∎
To finish the proof of Proposition 1.3 it suffices to study the case when $A_{\Gamma}$ is a free product of two free abelian groups.
Lemma 4.8.
Let $A_{\Gamma}=\mathbb{Z}^{n}\ast\mathbb{Z}^{m}$. Then, all $\ell^{2}$-Betti numbers of $\operatorname{Out}(A_{\Gamma})$ vanish except of two cases:
$\beta_{1}^{(2)}(\operatorname{Out}(F_{2}))=\frac{1}{24},\mathrm{and}~\beta_{2}^{(2)}(\operatorname{Out}(\mathbb{Z}^{2}\ast\mathbb{Z}^{2}))=\frac{1}{2^{11}3^{2}}$
Proof.
If $n=m=1$ then $A_{\Gamma}\cong F_{2}$ and it is well-known that $\operatorname{Out}(F_{2})=\mathrm{GL}_{2}(\mathbb{Z})$. In this case, the $\ell^{2}$-Betti numbers follow by Theorem 2.5. For all other cases, observe that $\operatorname{Out}(A_{\Gamma})$ contains no partial conjugations, so by the short exact sequence 1, we have $\mathrm{SOut}^{0}(A_{\Gamma})\cong Q$. If $n=1$ and $m=2$, or vice versa, $E(\Lambda_{\Gamma})\neq\mathrm{loop}(\Lambda_{\Gamma})$ and by Lemma 4.1.1 all $\ell^{2}$-Betti numbers vanish. If $n,m\geq 2$, then by Lemma 4.1.2 we have that all $\ell^{2}$-Betti numbers of $Q$ vanish except when $n=m=2$. In this case $\beta_{2}^{(2)}(\mathrm{SOut}^{0}(A_{\Gamma}))=\frac{1}{12^{2}}$. Since $\left|\operatorname{Out}(A_{\Gamma}):\mathrm{SOut}^{0}(A_{\Gamma})\right|=2^{7}$, it follows that $\beta_{2}^{(2)}(\operatorname{Out}(\mathbb{Z}^{2}\ast\mathbb{Z}^{2}))=\frac{1}{2^{7}}\frac{1}{12^{2}}=\frac{1}{2^{11}3^{2}}$. ∎
@@ 4.4 Higher dimensional $\ell^{2}$-Betti numbers
In general, computing the $\ell^{2}$-Betti numbers of $\operatorname{Aut}(A_{\Gamma})$ and $\operatorname{Out}(A_{\Gamma})$ for RAAGs based on a connected graph $\Gamma$ is a difficult problem. However, for several families of RAAGs, results in the literature imply that all $\ell^{2}$-Betti numbers of $\operatorname{Out}(A_{\Gamma})$ vanish.
@@ 5.1 First $\ell^{2}$-Betti number
This graph defines a RAAG $A_{\Gamma}$ with no transvections and no non-trivial graph automorphisms. In [Apendix A, [34]], Wiedmer proves using Theorem 2.26 that $\mathrm{PSO}(A_{\Gamma})\cong F_{2}$. Hence:
@@ 5.3 Algebraic fibring
Let $A_{\Gamma}$ be the RAAG from Example 5.1.1. By Theorem 3.19.2 and Proposition 3.17, the group $\operatorname{Aut}(A_{\Gamma}\times A_{\Gamma})$ virtually algebraically fibres.
@@
If $\Gamma$ contains a leaf and has trivial centre then $A_{\Gamma}$ is directly indecomposable. Hence, if $\Gamma_{1}$ and $\Gamma_{2}$ are two graphs each with a leaf and trivial centre then, by Theorem 3.19.2 and Proposition 3.17, $\operatorname{Aut}(A_{\Gamma_{1}}\times A_{\Gamma_{2}})$ virtually algebraically fibres.
@@
By Example 5.9 it follows that $\operatorname{Aut}(A_{\Gamma}\times A_{\Gamma})$ virtually algebraically fibres. Furthermore, the outer automorphism $\operatorname{Out}(A_{\Gamma})$ satisfies condition $3^{\prime}$ of Theorem 3.19, and is therefore virtually indicable. Since $A_{\Gamma}$ is both centreless and directly indecomposable, Proposition 3.17 implies that $\operatorname{Out}(A_{\Gamma}\times A_{\Gamma})$ virtually algebraically fibres. By Proposition 2.21 the group $\operatorname{Out}(A_{\Gamma})^{2}$ is a finite index subgroup of $\operatorname{Out}(A_{\Gamma}\times A_{\Gamma})$. As a result, by applying Theorem 4.9.6, we conclude that all $\ell^{2}$-Betti numbers of $\operatorname{Out}(A_{\Gamma})$, and therefore of $\operatorname{Out}(A_{\Gamma}\times A_{\Gamma})$, vanish.

### v3

@@ 3.2 Transvection subgroup of $\operatorname{Out}(A_{\Gamma})$
We define a directed graph $\Lambda_{\Gamma}$, called transvection graph, with vertex set $\{V_{1},\dots,V_{N}\}$ and such that there is a directed edge from $V_{i}$ to $V_{j}$ if $v\leq w$ for some $v\in V_{i}$ and $w\in V_{j}$. There is a loop from $V_{i}$ to itself if and only if $\left|V_{i}\right|\geq 2$. In this setting, the image of a transvection $\tau_{v_{i}}^{v_{j}}$ under $\rho$ is the elementary matrix $\mathrm{E}_{i,j}$ with $1$’s in the diagonal and in the $(i,j)$-entry and zeroes elsewhere. Thus, the matrices in $Q$ are block lower triangular, with diagonal blocks in $\mathrm{SL}_{\left|V_{i}\right|}(\mathbb{Z})$ and non-trivial off-diagonal blocks $(i,j)$ permitted precisely when there is an edge $V_{i}\to V_{j}$ in $\Lambda_{\Gamma}$.
@@
    There exist two singleton equivalence classes $u]=\{u\}$ and $v]=\{v\}$ with $u\leq v$, such that there is no vertex $w\neq u,v$ satisfying $u\leq w\leq v$.
@@
If property (P2) holds then there exist $v_{i}\leq v_{j}$ such that there is no $v_{r}\neq v_{i},v_{j}$ with $v_{i}\leq v_{r}\leq v_{j}$. Moreover, $v_{i}]=\{v_{i}\}$ and $v_{j}]=\{v_{j}\}$. Consider the set:
@@
and define the map $f:Q\to\mathbb{Z}$ by $f(E_{i,j})=1$ and $f(E_{k,l})=0$ if $E_{k,l}\in\mathcal{E}$. This is a well-defined group homomorphism because it preserves all relators of $Q$ from Theorem 3.10. We claim that $\ker(f)=\langle\mathcal{E}\rangle$. As $\ker(f)$ is normally generated by $\mathcal{E}$, it is enough to prove that for every $E_{k,l}\in\mathcal{E}$ there exists $R\in\langle\mathcal{E}\rangle$ such that $E_{i,j}E_{k,l}=RE_{i,j}$. If $i\neq l$ and $j\neq k$ then $E_{i,j}$ and $E_{k,l}$ commute so the claim follows. If $i=l$ and $j\neq k$ then $E_{k,j}\in\mathcal{E}$ and $E_{i,j}E_{k,i}=E_{k,i}E_{k,j}E_{i,j}$ and if $j=k$ and $i\neq l$ then $E_{i,l}\in\mathcal{E}$ and $E_{i,j}E_{j,l}=E_{j,l}(E_{i,l})^{-1}E_{i,j}$. We deduce that $\ker(f)$ is finitely generated.
@@
Remark 3.16.
A group $G$ is said to be indicable if it admits a non-trivial homomorphism $\varphi:G\to\mathbb{Z}$. The direct product of indicable groups is algebraically fibred (cf. [16, Theorem 1]), and if $G$ and $H$ are directly indecomposable centerless groups, then $\operatorname{Aut}(G)\times\operatorname{Aut}(H)$ is a finite-index subgroup of $\operatorname{Aut}(G\times H)$ by Proposition 2.22. Therefore, using the results on virtual indicability of automorphisms of RAAGs (cf. [2, 28]), one can construct examples of groups that virtually fibre. We will illustrate this method with explicit examples in Section 5.3.
@@ 4.1 Transvection subgroup of $\operatorname{Out}(A_{\Gamma})$
  • If there are partial conjugations but no transvections, then $\mathrm{PSO}(A_{\Gamma})$ is a finite index subgroup of $\operatorname{Out}(A_{\Gamma})$. By Proposition 3.7, we have that $\mathrm{PSO}(A_{\Gamma})$ virtually fibres if and only if $\mathrm{PSO}(A_{\Gamma})$ is not isomorphic to a RAAG $A_{\Theta}$ with $\Theta$ disconnected. Since virtual fibring implies vanishing of the first $\ell^{2}$-Betti number, it follows that $\beta_{1}^{(2)}(\operatorname{Out}(A_{\Gamma}))>0$ if and only if $\mathrm{PSO}(A_{\Gamma})$ is isomorphic to a RAAG $A_{\Theta}$ with $\Theta$ a disconnected graph (c.f. Corollary 2.12).∎
@@ 4.3 Higher dimensional $\ell^{2}$-Betti numbers
In general, computing the $\ell^{2}$-Betti numbers of $\operatorname{Aut}(A_{\Gamma})$ and $\operatorname{Out}(A_{\Gamma})$ is a difficult problem. However, for several families of RAAGs, results in the literature imply that all $\ell^{2}$-Betti numbers of $\operatorname{Out}(A_{\Gamma})$ vanish.
@@ 5.1 First $\ell^{2}$-Betti number
This graph defines a RAAG $A_{\Gamma}$ with no transvections and no non-trivial graph automorphisms. In [Appendix A, [33]], Wiedmer proves using Theorem 2.26 that $\mathrm{PSO}(A_{\Gamma})\cong F_{2}$. Hence:
@@ 5.3 Algebraic fibring
Let $A_{\Gamma}$ be the RAAG from Example 5.1. By [2, Theorem 1.6], $\operatorname{Aut}(A_{\Gamma})$ is virtually indicable. Therefore, $\operatorname{Aut}(A_{\Gamma}\times A_{\Gamma})$ virtually algebraically fibres by Remark 3.16.
@@
If $\Gamma$ contains a leaf and has trivial centre, then $A_{\Gamma}$ is directly indecomposable. Hence, if $\Gamma_{1}$ and $\Gamma_{2}$ are two graphs each with a leaf and trivial centre, then by [2, Theorem 1.6] the groups $\operatorname{Aut}(A_{\Gamma_{1}})$ and $\operatorname{Aut}(A_{\Gamma_{2}})$ are virtually indicable. Therefore, $\operatorname{Aut}(A_{\Gamma_{1}}\times A_{\Gamma_{2}})$ virtually algebraically fibres by Remark 3.16.
@@
By Example 5.9 it follows that $\operatorname{Aut}(A_{\Gamma}\times A_{\Gamma})$ virtually algebraically fibres. Furthermore, the outer automorphism group $\operatorname{Out}(A_{\Gamma})$ is virtually indicable by [2, Theorem 1.6]. Therefore, $\operatorname{Out}(A_{\Gamma}\times A_{\Gamma})$ virtually algebraically fibres by Remark 3.16. By Proposition 2.22, the group $\operatorname{Out}(A_{\Gamma})^{2}$ is a finite-index subgroup of $\operatorname{Out}(A_{\Gamma}\times A_{\Gamma})$. As a result, applying Theorem 4.4.6, we conclude that all $\ell^{2}$-Betti numbers of $\operatorname{Out}(A_{\Gamma})$, and hence of $\operatorname{Out}(A_{\Gamma}\times A_{\Gamma})$, vanish.